OpenStudy (anonymous):

I need help w/this problem, please ∛24a⁴b²x∛40a²b (Everything is under the radical on both sides)

OpenStudy (saifoo.khan):

Ok, so u want the answer?

OpenStudy (saifoo.khan):

\[\Large \sqrt[3]{960a^6b^3}\] \[\Large \sqrt[3]{64 \times15 a^6b^3}\]

OpenStudy (saifoo.khan):

That can we written as, \[\Large 4 a^2b \sqrt{15}\]

OpenStudy (anonymous):

You are so sweet! I swear I just e-mailed you (I used the e-mail in your profile) I really appreciate any help!!

OpenStudy (saifoo.khan):

Thanks! ^_^ I just got your email. replying in a sec.

OpenStudy (saifoo.khan):

Replied. :D

OpenStudy (anonymous):

You Rock! Thanks so much!

OpenStudy (saifoo.khan):

Lol, welcome. Did u got my email yet?

OpenStudy (anonymous):

Yes, just sent you a reply

OpenStudy (saifoo.khan):

replied. :D

OpenStudy (anonymous):

Sent you a reply!

OpenStudy (saifoo.khan):

Same here.

OpenStudy (saifoo.khan):

Hey! im back. :D

OpenStudy (anonymous):

Hi there

OpenStudy (saifoo.khan):

Check your email. :)

OpenStudy (anonymous):

I got it! I can't thank you enough!! You are the best!

OpenStudy (saifoo.khan):

Hehe. You welcome! :) Have a good night.

OpenStudy (anonymous):

you too!

OpenStudy (saifoo.khan):

^_^